OpenStudy (anonymous):

Janet's gross pay is $750, from which $31 is deducted for OASDI, $10 for Medicare, and $14 for income tax. What is her net pay?

OpenStudy (denonakavro):

Just subtract 31+10+14 from 750 so 750-(31+10+14)= 750-55=695